### v2

@@ 4.2 Video
Video processing is one of the fundamental directions in computer vision. The primary goal of video processing is to effectively master spatio-temporal representations across long contexts. Mamba excels in this domain with its selective state space model, which achieves a balance between maintaining linear complexity and enabling effective long-term dynamic modeling. This innovative approach has been widely adopted in diverse video analysis tasks such as video understanding (Li et al, 2024c) and video generation (Hu et al, 2024).
@@
Table 7: The Mamba-based methods in point cloud vision tasks. The abbreviations of Block here are VSS: Vision State Space Module, Vim: Vision Mamba, *: modification. The abbreviations of Task here are CLS: Classification, DET: Detection, SEG: Segmentation
Methods Block Data Task Code
@@
PointMamba (Liang et al, 2024) VSS* 3D Point Cloud Data Object CLS, Part SEG
PCM (Zhang et al, 2024d) Vim* 3D Point Cloud Data Object CLS, Part SEG, Semantic SEG
PointMamba (Liu et al, 2024d) Vim* 3D Point Cloud Data Object CLS, Semantic SEG
@@
Mamba3D (Han et al, 2024) Vim* 3D Point Cloud Data Object CLS, Part SEG
PoinTramba (Wang et al, 2024h) VSS 3D Point Cloud Data Object CLS, Part SEG
@@
LCM (Zha et al, 2024) Mamba 3D Point Cloud Data Object CLS, Object DET, Part SEG
3DMambaIPF (Zhou et al, 2024b) Mamba 3D Point Cloud Data Point Cloud Filtering
3DMambaC (Li et al, 2024g) Mamba* 3D Point Cloud Data Point Cloud Completion
EventMamba (Ren et al, 2024a) Mamba* 3D Event Cloud Data Action Recognition, Camera Pose Relocalization, Eye-tracking Regression
OverlapMamba (Xiang et al, 2024) Vim* LiDAR Data Place Recognition
4.2.1 Video Understanding
Mamba-ND (Li et al, 2024d) extended the Selective State Model to higher dimensions, unraveling the input data across various dimensions in row-major order. In addition to the 2D image tasks previously discussed, it also explored mamba’s effectiveness in video action recognition tasks. VideoMamba(Li et al, 2024c) extended Vim’s original 2D scan into different bidirectional 3D spatiotemporal scans to build a purely SSM-based model for video understanding. It showcased the scalability of the Mamba backbone in the visual domain through the use of self-distillation pretraining techniques. VideoMamba set new benchmarks for video understanding across short-term, long-term, and multi-modal video tasks, providing a scalable and efficient solution for comprehensive analysis of video content. Different from VideoMamba, VideoMamba Suite(Chen et al, 2024a) did not introduce a novel method but rather focused on assessing whether Mamba can serve as a viable alternative to Transformers in the video understanding domain. It leveraged Mamba in four distinct roles: temporal models, temporal modules, multi-modal interaction models, and space-time sequence models to reveal Mamba’s effectiveness in video understanding and to emphasize its key characteristics. VideoMamba Suite demonstrated Mamba’s ability to efficiently handle complex spatial-temporal dynamics, showcasing both superior performance and favorable efficiency-performance trade-offs. RhythmMamba (Zou et al, 2024a) designed multi-temporal Mamba as well as frequency domain feed-forward channel interacting to enable Mamba to capture the quasi-periodic patterns of remote photoplethysmography (rPPG). This study also explored the performance of RhythmMamba under long-time series inputs, demonstrating its robust performance over arbitrary length videos. Simba (Chaudhuri and Bhattacharya, 2024) leveraged Mamba within a novel encoder-decoder architecture with a Shift-GCN backbone to address the challenge of efficiently modeling long sequences inherent in skeleton action recognition tasks. In (Chen et al, 2024b), the authors curated the first AI-generated video detection dataset, GenVideo, and introduced the DeMamba module to enhance detectors for this task. MAMBA4D (Liu et al, 2024b), the first Mamba-based 4D backbone, was developed for generic point cloud video understanding.
Vivim (Yang et al, 2024e) introduced a Mamba-based framework for medical video object segmentation, featuring a novel temporal mamba block with spatiotemporal selective scanning.
4.2.2 Video Generation
SSMDiff (Oshima et al, 2024) replaced the attention layers in diffusion models with bidirectional SSMs to facilitate the generation of long video sequences. Zigma (Hu et al, 2024) developed several spatially continuous, space-filling 2D scanning schemes for image generation. Additionally, it factorized the 3D Mamba model into 2D and 1D Zigzag Mamba to enhance the integration of spatial and temporal information in video generation. Matten (Gao et al, 2024) investigated a latent diffusion model for video generation. It explored four model variants with different combinations of Mamba and attention mechanisms, revealing that integrating attention mechanisms for capturing local spatiotemporal details, and Mamba architectures for capturing global information, yields optimal performance.
4.3 Point Cloud
Point cloud is a fundamental 3D representation, which provides continuous spatial position information with 3D coordinates. The intrinsic disorder and irregularity natures of point cloud have been a challenge in 3D vision. Inspired by the linear complexity and global modeling capabilities of Mamba, several general SSM-based backbones (Liang et al, 2024; Zhang et al, 2024d; Liu et al, 2024d; Zhou et al, 2024b; Li et al, 2024g; Han et al, 2024) have been investigated in the field of point cloud processing.
In common tasks such as classification and segmentation, some methods primarily explore logical geometric scanning orders to convert point clouds into 1D point sequences. These methods include reordering strategy (Liang et al, 2024), consistent traverse serialization (Zhang et al, 2024d), and octree-based ordering mechanism (Liu et al, 2024d). Mamba3D (Han et al, 2024) employed a bidirectional SSM with channel flipping, and introduced Local Norm Pooling (LNP) block to extract local geometric features. PoinTramba (Wang et al, 2024h) is a hybrid Transformer-Mamba framework and introduced a bi-directional importance-aware ordering strategy. LCM (Zha et al, 2024) applied the Mamba architecture to Masked Point Modeling (MPM) pre-training, and proposed a locally constrained compact point cloud model.
In addition, 3DMambaIPF (Zhou et al, 2024b) was developed for iterative filtering to remove the noise from the point cloud. 3DMambaC (Li et al, 2024g) was designed to transform an initially incomplete and low-quality point cloud into a complete and high-fidelity version. EventMamba (Ren et al, 2024a) introduced a Mamba-based point cloud framework for event camera classification and regression tasks. OverlapMamba (Xiang et al, 2024) first transforms input range views into feature sequences and then processes these sequences using bidirectional Mamba for place recognition.
@@ 4.3 Point Cloud
Table 8: The Mamba-based methods in Multi-modal Vision Tasks. The abbreviations of Block here are VSS: Vision State Space Module, Vim: Vision Mamba, *: modification. The abbreviations of Data here are CT: Computed Tomography, PAN: Panchromatic Images, LRHS: Low-resolution Hyper-spectral Images, LRMS: Low-resolution Multi-spectral Images, MRI: Magnetic Resonance Imaging, PET: Positron Emission Tomography, SPECT: Single-photon Emission Computed Tomography. The abbreviations of Task here are CLS: Classification, DET: Detection, IVF: Infrared-visible Image Fusion, MIF: Medical Image Fusion, MCR: MRI-CT Registration, RSIP: Remote Sensing Images Pansharpening, SEG: Segmentation, SU: Scene Understanding, VQA: Visual Question Answering
       Methods        Block        Data        Task        Code
@@
       MambaDFuse (Li et al, 2024h)        Mamba*         RGB Images & Infrared Images, MRI & CT/PET/SPECT         IVF, MIF, Object DET
       FusionMamba (Xie et al, 2024b)        VSS*         RGB Images & Infrared Images, MRI & CT/PET/SPECT        IVF, MIF
       S4Fusion (Ma et al, 2024b)        VSS        RGB Images & Infrared Images        IVF
       Fusion-Mamba (Dong et al, 2024)        VSS*        RGB Images & Infrared Images        Object DET
       CFMW (Li et al, 2024b)        Mamba*        RGB Images & Infrared Images        Object DET
       Sigma (Wan et al, 2024)        VSS*        RGB Images & Thermal/Depth Images        Semantic SEG
       MambaMorph (Guo et al, 2024b)        Mamba        MRI & CT        MCR
       VMambaMorph (Wang et al, 2024i)        VSS        MRI & CT        MCR
       I2I-Mamba (Atli et al, 2024)        VSS*        Multi-Contrast MRI, MRI & CT        Image Generation
       PanMamba (He et al, 2024b)        VSS*        PAN & LRMS        RSIP
       FusionMamba (Peng et al, 2024)        VSS*        PAN & LRMS/LRHS, RGB & LRHS        RSIP
       LE-Mamba (Cao et al, 2024)        VSS*        PAN & LRMS/LRHS, RGB & LRHS        RSIP
@@
       InstructGIE (Meng et al, 2024)        VSS        Natural Images & Text        Image Editing
       VLMamba (Qiao et al, 2024)        VSS*        Natural Images & Text        VQA, Reasoning, SU
       Meteor (Lee et al, 2024)        Mamba        Natural Images & Text         VQA, Reasoning, SU
       Cobra (Zhao et al, 2024a)        Mamba        Natural Images & Text        VQA, Reasoning
       ReMamber (Yang et al, 2024d)        VSS*        Natural Images & Text        Referring Image SEG
       CLIP-Mamba (Huang et al, 2024f)        VSS/Mamba*        Natural Images & Text         Zero-shot CLS, Out of Distribution CLS
       SurvMamba (Chen et al, 2024g)        Vim*        Whole Slide Images & Gene        Survival Prediction
       RSCaMa (Liu et al, 2024a)        Vim*        Remote Sensing Images & Text        Change Captioning
       SpikeMba (Li et al, 2024e)        Vim*        Natural Videos & Text        Temporal Video Grounding
       BroadMamba (Shou et al, 2024)        Vim*        Audio & Video & Text        Emotion Recognition
       Coupled Mamba (Li et al, 2024f)        Mamba*        Audio & Video & Text        Sentiment Analysis
       Mamba-FETrack (Huang et al, 2024a)        Vim*        Video & Event Data        Frame-event Tracking
       MotionMamba (Zhang et al, 2024e)        Mamba*        Motion Sequence & Text        Motion Generation
       TM-Mamba (Wang et al, 2024f)        Vim*        Motion Sequence & Text        Human Motion Grounding
4.4 Multi-modal
@@
4.4.1 Homogeneous Multi-modal
@@
Homogeneous multi-modal paradigm refers to the tasks where the input data consists of multiple modalities sharing similar data type, including MRI-CT registration (Guo et al, 2024b; Wang et al, 2024i), medical image fusion (Li et al, 2024h; Xie et al, 2024b), medical image generation (Atli et al, 2024), remote sensing images pansharpening (Cao et al, 2024; He et al, 2024b; Peng et al, 2024), infrared-visible image fusion (Li et al, 2024h; Ma et al, 2024b; Xie et al, 2024b), RGB-infrared image object detection (Li et al, 2024b; Dong et al, 2024; Li et al, 2024h) and RGB-thermal/depth image semantic segmentation (Wan et al, 2024).
For the infrared-visible image fusion tasks in the natural domain, several methods (Li et al, 2024h; Ma et al, 2024b; Xie et al, 2024b) employed Mamba block to extract and integrate feature representations of RGB images and infrared images. Specifically, MambaDFuse (Li et al, 2024h) proposed a Mamba-based Dual-phase Fusion model, including dual-level single-modal feature extraction, dual-phase multi-modal feature fusion, and fused image reconstruction. FusionMamba (Xie et al, 2024b) proposed a dynamic feature enhancement method based on Mamba, which integrates VSS block with dynamic convolution and channel attention. S4Fusion (Ma et al, 2024b) proposed the Saliency-aware Selective State Space Fusion model, which can simultaneously focus on global spatial information from both modalities while facilitating their interaction. Meanwhile, Fusion-Mamba (Dong et al, 2024), MambaDFuse (Li et al, 2024h) and CFMW (Li et al, 2024b) aimed to utilize cross-modality complementary information from RGB images and infrared images to effectively improve object detection performance. Fusion-Mamba (Dong et al, 2024) introduced a Fusion-Mamba block designed to map cross-modal features into a hidden state space for interaction, thereby reducing disparities between these features and enhancing the consistency of the fused representations. CFMW (Li et al, 2024b) introduced a new task of handling adverse weather conditions in visible-infrared object detection, and designed Cross-Modality Fusion Mamba modules to enhance multi-modal integration. Meanwhile, Sigma (Wan et al, 2024) introduced a Mamba network for multi-modal semantic segmentation, which includes a siamese feature extractor, a feature fusion module, and an upsampling decoder.
In the medical domain, MambaMorph (Guo et al, 2024b) and VMambaMorph (Wang et al, 2024i) focused on the MRI-CT registration task and respectively integrated Mamba block and Visual State Space (VSS) block into the U-shaped encoder-decoder architectures for efficient long-range correspondence modeling. I2I-Mamba (Atli et al, 2024) was designed to enhance multi-modal image generation tasks by integrating channel-mixed Mamba blocks within the bottleneck of a convolutional backbone. Meanwhile, both MambaDFuse (Li et al, 2024h) and FusionMamba (Xie et al, 2024b) also focused on medical image fusion to learn structural and anatomical information from structural images such as CT and MRI and metabolic and functional information from functional images such as PET and SPECT.
For the remote sensing images pansharpening tasks in the remote sensing domain, PanMamba (He et al, 2024b) proposed channel swapping Mamba and cross-modal Mamba, designed for efficient information exchange and fusion between low-resolution multi-spectral and high-resolution panchromatic images. FusionMamba (Peng et al, 2024) incorporated VSS blocks into two U-shaped networks for extracting spatial and spectral features, and proposed the FusionMamba block by expanding SSM to accommodate dual inputs. LE-Mamba (Cao et al, 2024) introduced the Local Enhancement Vision Mamba block, which partitions the input image into windows for processing, and a state sharing technique to enhance the interaction between spatial and spectral information.
4.4.2 Heterogeneous Multi-modal

### v3

@@ 4.2. Video
Video processing is a key area in computer vision, dealing with 3D video data that consists of sequences of 2D frames. Video data includes a temporal dimension in addition to the spatial dimensions, enabling the capture of temporal dynamics. The primary goal of video processing is to effectively master spatio-temporal representations across long contexts. Mamba excels in this domain with its selective state space model, which achieves a balance between maintaining linear complexity and enabling effective long-term dynamic modeling. To adapt Mamba to 3D video data, current approaches organize the 3D video tokens in the following ways.
Spatial-first scans spatial tokens by location and then stacks them frame by frame. The scanning techniques for spatial tokens within each frame are similar to those used for 2D images. Temporal-first arranges temporal tokens by frame and then stacks them along the spatial dimension. To improve the receptive field and mitigate Mamba’s position sensitivity, these spatiotemporal scanning paths can be directly reversed. Considering the sequential attribute of the temporal dimension, Spatial Reversal reverses the spatial tokens in each frame and maintains the order along the temporal axis. Temporal Reversal maintains the order of spatial tokens and reverses the temporal sequence.
Additionally, some approaches apply Mamba separately to spatial and temporal information, utilizing 2D Spatial scanning and 1D Temporal scanning. Other approaches focus solely on Spatial scanning, leveraging alternative techniques to handle temporal dynamics. The video data can also be regarded as 3D data. In this situation, scanning techniques like 3D zigzag scanning and 3D Hilbert scanning can be utilized.
These innovative approaches, illustrated in Fig. 7, have been widely adopted in diverse video analysis tasks such as video understanding (Li et al., 2024c) and video generation (Hu et al., 2024).
4.2.1. Video Understanding
Recent approaches in video understanding tasks have adopted Mamba as the core architecture due to its superior long-range modeling capability and computational efficiency for analyzing high-resolution long video data, with a primary focus on optimizing scanning techniques (Yang et al., 2024e; Li et al., 2024c; Park et al., 2024a; Chen et al., 2024c; Zou et al., 2024b; Xiao et al., 2024a; Jin et al., 2024; Tang et al., 2024b; Chen et al., 2024b; Liu et al., 2024c). Both VideoMamba (Li et al., 2024c) and VideoMamba (Park et al., 2024a) built a purely Mamba-based model by stacking identical bi-directional Mamba blocks and utilizing the BD Spatial-first scanning strategy. VideoMamba (Li et al., 2024c) conducted extensive experiments comparing Spatial-first scanning, Temporal-first scanning, and their combination. The results indicate that BD Spatial-first scanning achieves the best performance, highlighting VideoMamba’s strong scalability, sensitivity to short-term action recognition, superior long-term video understanding, and compatibility with other modalities. VideoMamba (Park et al., 2024a) explored the reverse of Spatial-first scanning, conducting experiments on direct backward scanning, Spatial Reversal, and Temporal Reversal. The results reaffirm that BD Spatial-first scanning yields the best performance. Similarly, the Mamba architecture is employed for 3D long-term modeling in the multi-object tracking paradigm. MambaTrack (Xiao et al., 2024a) leveraged Mamba to model the spatio-temporal dynamics of objects and predict their future motion. DLT (Jin et al., 2024) employed Mamba to process multiple frames and estimate the distances of pre-detected 2D objects in these frames from a monocular camera. VMRNN (Tang et al., 2024b) integrated the 2D visual Mamba block with LSTM for spatiotemporal forecasting.
4.2.2. Video Generation
Video generation approaches focus on capturing correlations across the spatiotemporal dimension (Wu et al., 2024c; Oshima et al., 2024; Hu et al., 2024; Gao et al., 2024; Mo and Tian, 2024; Zhang et al., 2024e; Dong et al., 2024b). RainMamba (Wu et al., 2024c) introduced a 3D Hilbert scanning mechanism, while SSMDiff (Oshima et al., 2024), Zigma (Hu et al., 2024), and MMD (Dong et al., 2024b) decoupled spatial and temporal modeling. VFIMamba (Zhang et al., 2024e) utilized Temporal-first scanning to enhance spatiotemporal adjacency for the frame interpolation task.
@@ 4.2.2. Video Generation
Table 8. The Mamba-based methods in point cloud vision tasks. The abbreviations of Task here are CLS: Classification, DET: Detection, SEG: Segmentation. Tok: Tokenization.
Methods Data Task Tok. Scan Strategy Venue Code
@@
PointMamba (Liang et al., 2024) 3D Point Cloud Data Object CLS, Part SEG FPS & KNN & Hilbert/Trans-Hilbert SD Raster NeurIPS24
@@
Mamba3D (Han et al., 2024) 3D Point Cloud Data Object CLS, Part SEG FPS & KNN Raster + Channel Flipping MM24
LCM (Zha et al., 2024) 3D Point Cloud Data Object CLS, Object DET, Part SEG FPS & KNN SD Raster NeurIPS24
VoxelMamba (Zhang et al., 2024b) 3D Point Cloud Data Object DET Voxelization Multi-scale BD Hilbert NeurIPS24
@@
MambaMOS (Zeng et al., 2024) 4D Point Cloud Data Object SEG Z/Hilbert SD Raster MM24
4.3. Point Cloud
Point clouds represent a fundamental 3D data structure, consisting of discrete points in space, each defined by a set of 3D coordinates. These points collectively describe the geometry of objects or scenes. However, the inherent sparse and irregular distribution of point clouds poses significant challenges for 3D vision tasks. Point clouds can be used directly as raw points or transformed into regular grids through voxelization for further processing. When using raw points, they must first undergo a serialization process during tokenization. Notably, the computational complexity of 3D point cloud analysis can be reduced by grouping points or voxels. A common approach for grouping raw points during tokenization involves sampling key points using Farthest Point Sampling (FPS), followed by selecting k-nearest neighbors (KNN) for each key point.
Inspired by the linear complexity and global modeling capabilities of Mamba, several general SSM-based backbones (Zubic et al., 2024; Liang et al., 2024; Zhang et al., 2024d; Liu et al., 2024h; Han et al., 2024; Wang et al., 2024c; Zha et al., 2024; Zhang et al., 2024b; Zhou et al., 2024c; Li et al., 2024i; Ren et al., 2024b; Xiang et al., 2024; Zeng et al., 2024; Liu et al., 2024c; Mo, 2024) have been explored in point cloud processing. Adaptations of 1D causal Mamba operations for 3D point clouds typically focus on scanning strategies and tokenization of raw points. For example, Z-order and Hilbert curves can be employed during scanning or tokenization to preserve spatial locality.
PointMamba (Liang et al., 2024) applied Hilbert and Trans-Hilbert curves to the key points obtained through FPS before performing KNN during tokenization. This study demonstrated that PointMamba, leveraging a simple Mamba block, outperforms state-of-the-art Transformer-based networks. Mamba3D (Han et al., 2024) identified that learning sequence order in point clouds can lead to unreliable and unstable pseudo-order dependencies. To mitigate this, channel flipping was introduced for the backward SSM. LCM (Zha et al., 2024) applied the Mamba architecture to masked point modeling for pre-training, and proposed a locally constrained compact point cloud model. VoxelMamba (Zhang et al., 2024b) introduced a Multi-scale BD Hilbert scanning strategy. MambaMOS (Zeng et al., 2024) employed Z-order and Hilbert curves to serialize 4D points and designed a new embedding layer to better capture spatial and temporal information.
@@ 4.3. Point Cloud
Table 9. The Mamba-based methods in Multi-modal Vision Tasks. The abbreviations of Task here are DET: Detection, SEG: Segmentation, SU: Scene Understanding, VQA: Visual Question Answering. The abbreviations of Block here are VSS: Vision State Space Module, *: modification.
Methods Data Task Fusion Strategy Venue Code
Paradigm Mamba Location Mamba Design
@@
MGMF (Wang et al., 2024k) RGB Images & Infrared Images Object DET Multi-branch Fusion Module M TGRS24
Sigma (Wan et al., 2025) RGB Images & Thermal/Depth Images Semantic SEG Multi-branch Backbone/Fusion Module/Decoder VSS/M/U/VSS* WACV25
@@
InstructGIE (Meng et al., 2024) Natural Images & Text Image Editing Diffusion Vision Encoder VSS ECCV24
Meteor (Lee et al., 2024) Natural Images & Text VQA, Reasoning, SU MLLM LLM Mamba NeurIPS24
RoboMamba (Liu et al., 2024d) Natural Images & Text Robot Reasoning and Manipulation MLLM LLM Mamba NeurIPS24
ReMamber (Yang et al., 2024c) Natural Images & Text Referring Image SEG Uni-branch Backbone U ECCV24
RSCaMa (Liu et al., 2024a) Remote Sensing Images & Text Change Captioning Multi-branch Fusion Module M/U GRSL24
CoupledMamba (Li et al., 2024k) Audio & Video & Text Sentiment Analysis Uni-branch Backbone M NeurIPS
MotionMamba (Zhang et al., 2024f) Motion Sequence & Text Motion Generation Diffusion Backbone CA ECCV24
MambaTalk (Xu et al., 2024a) Audio & Text & Motion Motion Generation VQ-VAE Fusion Module CA/U NeurIPS24
MambaGesture (Fu et al., 2024b) Audio & Text & Style & Emotion Motion Generation Diffusion Backbone CA/A MM24
4.4. Multi-modal
Multi-modal tasks have become increasingly pivotal in the domain of computer vision, enabling the enrichment of visual understanding through the integration of diverse information sources. The fundamental objective of multi-modal tasks is to learn valuable potential feature representations from a multitude of modalities, such as textual captions and visual images, RGB images with supplementary components like depth or thermal images, and various forms of medical imaging data. Nevertheless, a significant challenge in attaining multi-modal objectives lies in the effective capture and amalgamation of the information across diverse modalities.
Early multi-modal approaches employ the Multi-branch paradigm, where each branch processes distinct modalities in parallel, followed by the fusion of features to create a unified representation. With the advent of Transformer architectures, a new Uni-branch paradigm has emerged, which concatenates various modalities and processes them within a single branch, simplifying the integration process and achieving greater efficiency. To harness the potential of large language models (LLMs), the Multi-modal Large Language Model (MLLM) paradigm utilizes LLMs as task decoders, facilitating enhanced multi-modal understanding. Additionally, established paradigms such as Diffusion (Ho et al., 2020) and VQ-VAE (van den Oord et al., 2017) are widely adopted for multi-modal generation tasks, leveraging their robust generative capabilities.
Recently, Mamba has been adapted for a variety of multi-modal tasks due to its computational efficiency and flexibility in feature fusion, demonstrating impressive performance across different applications. Firstly, various modules within existing paradigms can be replaced with Mamba-based architectures to leverage its global receptive capabilities and linear computational complexity, such as backbones and vision encoders. Notably, the LLM component of MLLM paradigms can be substituted with Mamba to capitalize on its advantages. Secondly, leveraging Mamba’s selective SSM mechanism, novel Mamba designs can enhance feature fusion, enabling flexible interactions between features from different modalities. Some fusion approaches employ distinct SSM mechanisms for each modality, integrating information from other modalities (individual, combined, or differences between modalities) into the SSM framework (M). Some fusion approaches concatenate information from different modalities, leveraging the SSM mechanism to fuse them through advanced scanning techniques (U). Additionally, mechanisms such as Cross Attention (CA) and Attention (A) can be incorporated into the Mamba-based architecture to facilitate effective feature fusion.
@@ 4.4. Multi-modal
4.4.1. Homogeneous Multi-modal
@@
Homogeneous multi-modal applications refer to the tasks where the input data consists of multiple modalities sharing similar data type, including MRI-CT registration (Guo et al., 2024c; Wang et al., 2024n), medical image fusion (Li et al., 2024f; Xie et al., 2024a), medical image generation (Atli et al., 2024), remote sensing images pansharpening (He et al., 2024b; Peng et al., 2024; Cao et al., 2024), infrared-visible image fusion (Li et al., 2024f; Xie et al., 2024a; Ma et al., 2024c), RGB-infrared image object detection (Dong et al., 2024d; Li et al., 2024b, f; Wang et al., 2024k) and RGB-thermal/depth image semantic segmentation (Wan et al., 2025).
For RGB-infrared object detection, the multi-branch paradigm MGMF (Wang et al., 2024k) introduced the Mamba-based fusion module, while utilizing the common feature extraction backbone and detection head. In particular, this fusion module maps the summed features from the two modalities into a hidden state space and employs gating mechanisms to facilitate state transitions, enabling deep cross-modal fusion of the representations. Sigma (Wan et al., 2025) is another multi-branch paradigm designed for multi-modal semantic segmentation, employing a fully Mamba-based architecture across all components. Specifically, VMamba (Liu et al., 2024e) blocks are used in the feature extraction backbone. For the fusion module, cross-modal multiplication and feature concatenation mechanisms are integrated within the SSM. Furthermore, channel attention is incorporated into the SSM to enhance the decoder’s capability in modeling channel-wise interactions.
4.4.2. Heterogeneous Multi-modal
Heterogeneous multi-modal applications refer to tasks where the input data consists of multiple modalities with different data types (Meng et al., 2024; Qiao et al., 2024; Lee et al., 2024; Zhao et al., 2024b; Liu et al., 2024d; Yang et al., 2024c; Huang et al., 2024e; Chen et al., 2024g; Liu et al., 2024a; Li et al., 2024a; Shou et al., 2024; Li et al., 2024k; Huang et al., 2024f; Zhang et al., 2024f; Fu et al., 2024b; Wang et al., 2024f; Sun et al., 2024).